# No [CII] or dust detection in two Little Red Dots at

### v1

@@ 2.2 NOEMA observations
The NOEMA observations of ID9094 and ID2756 were carried out in the summer of 2023 using the PolyFiX correlator (Project ID: S23CY, PIs: D. Elbaz & M. Xiao). Conducted with array configuration D in band 3, the observations targeted the [C ii] emission line, and the 1.3 mm dust continuum, with representative frequencies of 232.052 GHz for ID9094 and 236.576 GHz for ID2756. The total on-source time was 6 hours and 4.5 hours for ID2756 and ID9094, respectively.
All the calibrations and the creation of table were performed using CLIC package from the IRAM GILDAS software333https://www.iram.fr/IRAMFR/GILDAS/ with the support of IRAM astronomers. We then performed further analysis with version 6.5.5 of the Common Astronomy Software Application package (CASA; McMullin et al., 2007).
For the [C ii] emission line, imaging was carried out using the tclean task with 0.4′′ pixels and a channel width of 50 km s-1 with a natural weighting. The natural weighting provides the best point-source sensitivity, which is optimal for source detection. The resulting data cube has a synthesized beam size of full width at half maximum (FWHM) 1838 1472 (9.8 kpc 7.8 kpc in physical scale) and 2859 1854 ( 15.0 kpc 9.7 kpc) with a root mean square (rms) sensitivity of 0.45 and 0.84 mJy beam-1 per channel at the phase center, for ID9094 and ID2756, respectively.
We also created the observed 1.3 mm continuum maps using natural weighting and the same method as for the [C ii] emission line. For ID9094, the rms level is 30.4 Jy beam-1 in the map of 1725 1383 ( 9.2 kpc 7.4 kpc) angular resolution. For ID2756, the rms level is 50.3 Jy beam-1 in the map of 2712 1743 ( 14.3 kpc 9.2 kpc) angular resolution.
@@ 2.3 Multi-wavelength observations
Besides the JWST FRESCO and NOEMA observations, the two sources are also covered by a wide array of multi-wavelength observations. In Table 1, we summarize the multi-wavelength dataset for our two sources from X-ray to radio: () X-ray: Chandra 2 Ms (0.5–7.0keV, 0.5–2.0keV, and 2–7keV bands) images in the Chandra Deep Field-North field (CDF-N; Xue et al., 2016); () HST/ACS (F435W, F606W, F775W, F814W, F850LP) and HST/WFC3 (F105W, F125W, F140W, F160W) images from the Hubble Legacy Fields Program (HLF444https://archive.stsci.edu/prepds/hlf/; Whitaker et al., 2019); () JWST/NIRCam (F090W, F115W, F150W, F182M, F200W, F210M, F356W, F444W) from FRESCO (Oesch et al., 2023), CONGRESS (Egami et al., 2023), and JADES (Eisenstein et al., 2023; D’Eugenio et al., 2024). The two sources are not yet covered by any publicly available JWST/MIRI observations; () Spitzer/IRAC from the GREATS program (3.6 µm, 4.5 µm, 5.8 µm, 8 µm; Stefanon et al., 2021, new catalogs based on FRESCO sources are taken from Stefanon et al. in prep; priv. comm.); () Spitzer/MIPS, Herschel/PACS and SPIRE catalog is taken from the GOODS-Herschel survey (Elbaz et al., 2011); () JCMT/SCUBA-2 at 850 µm (Cowie et al., 2017); () NOEMA data from this study (see Sect. 2.2); () NIKA2 Cosmological Legacy Survey at 1.2 mm and 2.0 mm (Bing et al., 2023); and () VLA 1.5 GHz (Owen, 2018).
@@
We note that the photometric measurements from HST and JWST images are derived following the same procedure as outlined in Weibel et al. (2024). Briefly, we use SExtractor (Bertin & Arnouts, 1996) in dual image mode with an inverse-variance weighted stack of the F210M+F444W with FRESCO-only imaging, as the detection image. In this study, fluxes are measured in 016 radius circular apertures in images that are point spread function (PSF)-matched to the F444W band. Total fluxes are derived from the Kron AUTO aperture provided by SExtractor in the F444W band, in addition to a correction based on the encircled energy of the Kron aperture on the F444W PSF. Detailed descriptions of data reduction and photometric measurements are provided in Weibel et al. (2024) and Xiao et al. (2024).
2.4 Our sample: two “Little Red Dots” at
@@
Although there is no uniform definition of LRDs, we identify our two sources as LRDs because they satisfy both of two commonly used selection criteria: 1) an initial criterion – sources with broad Balmer emission lines ( km s-1; Matthee et al., 2024); and 2) a widely used criterion based only on photometric measurements – a combination of “v-shape” SED and compactness selection (e.g., Labbe et al., 2023; Kokorev et al., 2024; Greene et al., 2024).
@@
Fig. 1 presents the JWST/NIRCam images and grism spectra of ID9094 and ID2756. The RGB images show that both sources appear red with point-like morphology. The grism spectra show H and [O iii] emission lines. Both sources exhibit a broad H line. We fit the H line with both narrow and broad Gaussian components, assuming the narrow component has the same width as the [O iii] lines. The best-fit line properties are shown in Table 2. The broad component of H has a FWHM of km s-1 for ID9094 and km s-1 for ID2756, suggesting a possible Type I (broad emission line) AGN. Notably, ID2756 has also been identified as a strong Lyman- emitter, with a considerable equivalent width of 221 Å (Leonova et al. in prep.).
@@
We further compare our two sources to an alternative color and compactness selection criteria for LRDs, designed to identify compact, red objects with “v-shape” SEDs (e.g., Kokorev et al., 2024; Greene et al., 2024):
@@
Both sources are bright in F444W, with magnitudes of mag (ID9094) and mag (ID2756), consistent with the LRD selection. Since ID2756 lacks observations in F277W and F200W, while ID9094 is missing F277W, the corresponding fluxes used here are derived from the best-fit SED and the 16th–84th percentile range (as shown later in Fig. 4). The detailed comparison is shown in Fig. 2. ID9094 lies within the red region, perfectly matching the LRD selection. ID2756 exhibits a ”v-shaped” color and compact morphology, aligning with the LRD classification. Its F277W – F356W color is slightly bluer than the selection threshold but still within the 1 uncertainty. We therefore classify both sources as LRDs. Compared to the majority of sources from various JWST blank legacy fields (blue-shaded region in Fig. 2), obtained from the DAWN JWST Archive (DJA555https://dawn-cph.github.io/dja/imaging/v7/; Valentino et al., 2023), these two LRDs stand out as particularly exceptional, exhibiting redder colors and more compact morphologies.
@@
Altogether, we find that both sources are consistent with standard criteria for LRDs based on their morphology and “v-shape” SED. In addition, both sources show the broad H line, making them strong Type I AGN candidates.
@@ 3 Results
3.1 Non-detections in both [C ii] and 1.3 mm of two LRDs
We conduct NOEMA observations targeting the [C ii] emission line and 1.3 mm continuum, generating data cubes and 1.3 mm continuum maps (see Sect. 2.2). Using the data cube, we extract spectra at the positions of our sources with an aperture consistent with the NOEMA beam size. The resulting spectra, shown in Fig. 3, reveal no significant signal at the expected [C ii] frequency (). Assuming a [C ii] line width of 250 km s-1 – a typical value for high-redshift galaxies (Béthermin et al., 2020; Endsley et al., 2022) – we construct the velocity-integrated intensity map (moment-0). Fig. 3 overlays the contours from the [C ii] moment-0 and 1.3 mm maps on the JWST RGB image, showing no detections () at the source positions and in the surrounding regions. We therefore conclude that neither the [C ii] emission line nor the 1.3 mm dust continuum emission are detected for the two LRDs in the current NOEMA data.
We then determine 3 upper limits for both [C ii] and dust continuum. Using the rms of the 1.3 mm map (see Sect. 2.2) and assuming our sources are not resolved, we derive 3 upper limits of 91 Jy and 151 Jy for ID9094 and ID2756, respectively. In addition, the 3 upper limits for [C ii] are calculated as three times the pixel-by-pixel rms of the moment-0 map, which gives upper limits of 0.157 Jy km s-1 and 0.310 Jy km s-1 for ID9094 and ID2756, respectively. The corresponding 3 upper limits on the [C ii] emission-line luminosity are log for ID9094 and log for ID2756. Using , we further calculate the 3 upper limits on the molecular gas mass (Zanella et al., 2018) to be log for ID9094 and log for ID2756.
Intriguingly, when requesting the NOEMA observations, we assumed these two galaxies to be DSFGs. Given their stellar masses from Xiao et al. (2024), we estimated their total star formation rates (SFRtot) assuming that they lie on the star-forming main sequence (SFMS; Schreiber et al., 2015) and predicted their expected [C ii] fluxes using the –SFRtot relation for SFGs (Schaerer et al., 2020). Similarly, we predicted their 1.3 mm continuum fluxes by assuming an infrared SED shape consistent with typical dust-obscured galaxies (optically dark/faint galaxies; Xiao et al., 2023). Given the depth of our NOEMA observations, we expected ID9094 to be detected at in [C ii] and in the 1.3 mm continuum, while ID2756 should be detected in [C ii] at . However, the complete absence of [C ii] and 1.3 mm emission suggests that these two LRDs are not the dusty SFGs initially assumed. This result aligns with our findings in Sect. 2.4, where both sources exhibit a broad H emission line, being Type-I AGN candidates.
@@ 3.2 The hints of AGNs in two LRDs from FIR non-detections
In this section, we utilize the 3 upper limits on the [C ii] emission line and the 1.3 mm continuum to investigate the possible presence of AGN activity in the LRDs. While both sources have already been identified as strong Type-I AGN candidates based on their broad H emission lines, our primary goal here is to investigate whether the FIR non-detections can effectively rule out the DSFG scenario for LRDs, independently of spectral information. Here, we focus primarily on ID9094, which benefits from deeper NOEMA observations compared to ID2756, allowing for more robust constraints. We assess the AGN contribution through two independent approaches: 1) UV-to-FIR spectral energy distribution (SED) fitting analysis and 2) the –SFRtot empirical relation.
@@ 3.2.1 SED fitting analysis
We perform a UV-to-FIR SED fitting analysis using CIGALE (Boquien et al., 2019), which is based on an energy balance principle, fixing the redshift to . We assume a constant star formation history (SFH) and adopt Bruzual & Charlot (2003) stellar population models with the Calzetti et al. (2000) dust attenuation law. We adopt a broad metallicity range of 0.004 to 0.02, dust attenuation to the rest-frame V band () values from 0 to 6 magnitudes, and ionization parameters log of -2.0. The AGN component is modeled using the Fritz et al. (2006) template, with the AGN fraction (), i.e., the contribution of the AGN to the total IR luminosity, ranging from 0.01 to 0.8. The angle between the equatorial axis and the viewing angle at the line of sight () ranges from 40 to 90. With a fixed open angle of the AGN torus at 100 degrees, this range of refers to Type I AGN. We note that recent MIR observations (Wang et al., 2024a; Williams et al., 2024; Barro et al., 2024) suggest a lack of hot torus dust emission in LRDs, which may deviate from the Fritz et al. (2006) template adopted. However, with the absence of deep MIR observations, we anticipate little change in the quality of SED fitting on our sample by only changing the AGN torus emission in the templates. On the contrary, the inclusion of continuum from accretion disk by limiting to Type I AGN is critical, as we robustly detected broad lines in H.
@@
Given that we only have upper limits in the FIR, we adopt a simple FIR dust model from Casey (2012), using the default dust emissivity index () and MIR power-law slope (). During the fitting process, we find that when the dust temperature () is allowed to vary across a defined range, the best-fit is strongly influenced by the upper boundary of the input range. This effect arises due to the absence of deep MIR data, which prevents us from constraining the maximum . Consequently, we cannot determine a precise best-fit value for , but the deep NOEMA data allow us to place a lower limit.
@@
To refine our constraints, we adopt an optimized fitting approach that allows us to constrain the minimum allowed using the NOEMA 1.3 mm 3 upper limits. In this approach, we fix during the fitting and perform multiple fits with different values, ranging from 30 K to 150 K. In addition, we also correct the effect of the cosmic microwave background (CMB) as an observing background on photometry, following da Cunha et al. (2013). The CMB has the effect of raising the apparent dust temperature with increasing redshift, as the CMB temperature scales with redshift and provides an additional background source of radiation for dust heating. We assume , consistent with previous studies for galaxies (Sommovigo et al., 2022). In this way, we derive the intrinsic FIR photometry starting from MIPS 70m, which is used for SED fitting. The best-fit SEDs, with and without AGN components considered, are presented in Fig. 4.
@@
For ID9094, without including an AGN component (Fig.4-), we present three representative examples of best-fit SEDs with () K, () K, and () K. These correspond to results that the predicted flux density at 1.3 mm exceeds, matches, and falls below the observed 3 upper limits, respectively, under the energy balance assumption of CIGALE. At K – a typical value for star-forming UV-bright galaxies in the REBELS survey at (Sommovigo et al., 2022) – the predicted 1.3 mm flux exceeds the observed 3 upper limit by nearly a factor of three. This discrepancy suggests that, with our current NOEMA observations, we should have detected the source at a significance level of , which is clearly not the case. As increases, the peak of the FIR SED moves towards the shorter wavelength and the fit improves. At K (Fig.4-), the predicted 1.3 mm flux aligns with the observed 3 upper limit. However, given that the actual signal-to-noise ratio at 1.3 mm is well below 3, this implies that the real must be greater than 110 K.
@@
Therefore, without including an AGN component, we obtain a lower limit of K. At fixed K, as in Table 3, the best-fit IR luminosity is , yr-1 (we use this for the analysis in Sect. 3.2.2), and stellar mass of . Remarkably, even this lower limit is significantly higher than the typical K observed in REBELS galaxies at similar redshifts (Sommovigo et al., 2022). Similar findings have been reported in studies of 675 LRDs at , regardless of heating mechanism, whether AGN or star formation, where an average luminosity-weighted dust temperature of K is reported (Casey et al., 2024). Such an extreme dust temperature is challenging to explain solely through heating by star formation. In typical star-forming galaxies, the dust temperature is primarily driven by the ultraviolet radiation field of young stars, which is insufficient to reach the high values observed here.
@@
Based on the best-fit without AGN at fixed K, we derived the dust emitting region size is pc, based on a black-body assumption. A crucial parameter to consider is the stellar surface density, given by pc-2 (similar findings are reported in, e.g., Baggen et al., 2024). This is about one order of magnitude higher than the densest systems in the local universe (Hopkins et al., 2010) and SFGs at observed by JWST (Morishita et al., 2024; Schaerer et al., 2024), highlighting a general issue for LRD models that rely solely on stellar components. Expanding the star formation region would reduce , but it would also decrease below the observed values (e.g., De Rossi et al., 2018), leading to an inconsistency with our findings. All these results suggest that additional heating mechanisms, such as AGN-driven radiation, must play a significant role in explaining the elevated dust temperatures of ID9094.
| ID9094 | fixed minimum [K] | log ] | SFR [yr-1] | [] | |
| Without AGN | 110 | ||||
| With AGN | 60 |
@@
When the AGN component is included in the SED fitting (Fig. 4-), the results become more physically plausible. Similar to Fig.4-, we also show three representative examples of best-fit SEDs with () K, () K, and () K. These correspond to results that the predicted flux density at 1.3 mm exceeds, matches, and falls below the observed 3 upper limits, respectively. The best-fit results indicate the lower limit of K, which is consistent with AGN and/or quasar host galaxies at ( K; e.g., Walter et al., 2022; Tsukui et al., 2023; Tripodi et al., 2023; Decarli et al., 2023; Meyer et al., 2025). At fixed K, the AGN fraction is , indicating a scenario where both AGN and star formation contribute significantly to the FIR emission. The corresponding best-fit , yr-1, and (see Table 3).
@@
For ID2756, we perform SED fitting using the same strategy as for ID9094 (see Appendix A). However, the current 1.3 mm observations for ID2756 (rms level 50.3 Jy beam-1) are not sufficiently deep to constrain the SED effectively. The lack of sensitivity prevents us from distinguishing between the DSFG and AGN scenarios based solely on the SED fitting results. To achieve meaningful constraints, deeper FIR/submillimeter observations are required. Specifically, observations at least as deep as for ID9094 (rms level 30.4 Jy beam-1) are necessary to provide tighter limits on the 1.3 mm flux density and help refine the SED fitting.
In summary, our analysis highlights the role of deep FIR observations in distinguishing between DSFG and AGN-dominated scenarios for LRDs. For ID9094, the FIR SED fitting indicates that the best-fit results, including the high values, are inconsistent with typical SFGs and suggest the presence of an AGN. For ID2756, deeper FIR data are needed to drive comparable constraints. Despite the non-detection in the FIR, the energy balance imposed by the deep NOEMA upper limits provides meaningful constraints on the shape of the FIR SED. These results demonstrate the importance of FIR data in studying the nature of LRDs and ruling out purely star formation-driven scenarios.
3.2.2 –SFRtot relation

### v2

@@ 2.2 NOEMA observations
The NOEMA observations of ID9094 and ID2756 were carried out in the summer of 2023 using the PolyFiX correlator (Project ID: S23CY, PIs: D. Elbaz & M. Xiao). Conducted with array configuration D in band 3, the observations targeted the [Cāii] emission line, and the 1.3 mm dust continuum, with representative frequencies of 232.052 GHz for ID9094 and 236.576 GHz for ID2756. The total on-source time was 6 hours and 4.5 hours for ID2756 and ID9094, respectively.
All the calibrations and the creation of table were performed using CLIC package from the IRAM GILDAS software444https://www.iram.fr/IRAMFR/GILDAS/ with the support of IRAM astronomers. We then performed further analysis with version 6.5.5 of the Common Astronomy Software Application package (CASA; McMullin etĀ al., 2007).
For the [Cāii] emission line, imaging was carried out using the tclean task with 0.4ā²ā² pixels and a channel width of 50 km s-1 with a natural weighting. The natural weighting provides the best point-source sensitivity, which is optimal for source detection. The resulting data cube has a synthesized beam size of full width at half maximum (FWHM) 1838 1472 (9.8 kpc 7.8 kpc in physical scale) and 2859 1854 ( 15.0 kpc 9.7 kpc) with a root mean square (rms) sensitivity of 0.45 and 0.84 mJy beam-1 per channel at the phase center, for ID9094 and ID2756, respectively.
We also created the observed 1.3 mm continuum maps using natural weighting and the same method as for the [Cāii] emission line. For ID9094, the rms level is 30.4 Jy beam-1 in the map of 1725 1383 ( 9.2 kpc 7.4 kpc) angular resolution. For ID2756, the rms level is 50.3 Jy beam-1 in the map of 2712 1743 ( 14.3 kpc 9.2 kpc) angular resolution.
@@ 2.3 Multi-wavelength observations
Besides the JWST FRESCO and NOEMA observations, the two sources are also covered by a wide array of multi-wavelength observations. In Table 1, we summarize the multi-wavelength dataset for our two sources from X-ray to radio: () X-ray: Chandra 2 Ms (0.5ā7.0keV, 0.5ā2.0keV, and 2ā7keV bands) images in the Chandra Deep Field-North field (CDF-N; Xue etĀ al., 2016); () HST/ACS (F435W, F606W, F775W, F814W, F850LP) and HST/WFC3 (F105W, F125W, F140W, F160W) images from the Hubble Legacy Fields Program (HLF555https://archive.stsci.edu/prepds/hlf/; Whitaker etĀ al., 2019); () JWST/NIRCam (F090W, F115W, F150W, F182M, F200W, F210M, F356W, F444W) from FRESCO (Oesch etĀ al., 2023), CONGRESS (Egami etĀ al., 2023), and JADES (Eisenstein etĀ al., 2023; DāEugenio etĀ al., 2025). The two sources are not yet covered by any publicly available JWST/MIRI observations; () Spitzer/IRAC from the GREATS program (3.6 āµm, 4.5 āµm, 5.8 āµm, 8 āµm; Stefanon etĀ al., 2021, new catalogs based on FRESCO sources are taken from Stefanon et al. in prep; priv. comm.); () Spitzer/MIPS, Herschel/PACS and SPIRE catalog is taken from the GOODS-Herschel survey (Elbaz etĀ al., 2011); () JCMT/SCUBA-2 at 850 āµm (Cowie etĀ al., 2017); () NOEMA data from this study (see Sect. 2.2); () NIKA2 Cosmological Legacy Survey at 1.2 mm and 2.0 mm (Bing etĀ al., 2023); and () VLA 1.5 GHz (Owen, 2018).
@@
We note that the photometric measurements from HST and JWST images are derived following the same procedure as outlined in Weibel etĀ al. (2024). Briefly, we use SExtractor (Bertin & Arnouts, 1996) in dual image mode with an inverse-variance weighted stack of the F210M+F444W with FRESCO-only imaging, as the detection image. In this study, fluxes are measured in 016 radius circular apertures in images that are point spread function (PSF)-matched to the F444W band. Total fluxes are derived from the Kron AUTO aperture provided by SExtractor in the F444W band, in addition to a correction based on the encircled energy of the Kron aperture on the F444W PSF. Detailed descriptions of data reduction and photometric measurements are provided in Weibel etĀ al. (2024) and Xiao etĀ al. (2024).
2.4 Our sample: two āLittle Red Dotsā at
@@
Although there is no uniform definition of LRDs, we identify our two sources as LRDs because they satisfy both of two commonly used selection criteria: 1) an initial criterion ā sources with broad Balmer emission lines ( km s-1; Matthee etĀ al., 2024); and 2) a widely used criterion based only on photometric measurements ā a combination of āv-shapeā SED and compactness selection (e.g., Labbe etĀ al., 2025; Kokorev etĀ al., 2024; Greene etĀ al., 2024; Kocevski etĀ al., 2025).
@@
Fig.Ā 1 presents the JWST/NIRCam images and grism spectra of ID9094 and ID2756. The RGB images show that both sources appear red with point-like morphology. The grism spectra show H and [Oāiii] emission lines. Both sources exhibit a broad H line. We fit the H line with both narrow and broad Gaussian components, assuming the narrow component has the same width as the [Oāiii] lines. The best-fit line properties are shown in TableĀ 2. The broad component of H has a FWHM of km s-1 for ID9094 and km s-1 for ID2756, suggesting a possible Type I (broad emission line) AGN. Notably, ID2756 has also been identified as a strong Lyman- emitter, with a considerable equivalent width of 221 Ć Ā (Leonova et al. in prep.).
@@
We further compare our two sources to an alternative color and compactness selection criterion for LRDs, designed to identify compact, red objects with āv-shapeā SEDs (e.g., Kokorev etĀ al., 2024; Greene etĀ al., 2024):
@@
Both sources are bright in F444W, with magnitudes of mag (ID9094) and mag (ID2756), consistent with the LRD selection. Since ID2756 lacks observations in F277W and F200W, while ID9094 is missing F277W, the corresponding fluxes used here are derived from the best-fit SED and the 16thā84th percentile range (as shown later in Fig.Ā 4). The detailed comparison is shown in Fig.Ā 2. ID9094 lies within the red region, perfectly matching the LRD selection. ID2756 exhibits a āv-shapedā color and compact morphology, aligning with the LRD classification. Its F277W ā F356W color is slightly bluer than the selection threshold but still within the 1 uncertainty. We therefore classify both sources as LRDs. Compared to the majority of sources from various JWST blank legacy fields (blue-shaded region in Fig.Ā 2), obtained from the DAWN JWST Archive (DJA666https://dawn-cph.github.io/dja/imaging/v7/; Valentino etĀ al., 2023), these two LRDs stand out as particularly exceptional, exhibiting redder colors and more compact morphologies.
To further quantify the āv-shapeā SEDs, we also measured the UV and optical continuum slopes ( and ) using multi-band photometry, following the method of Kocevski etĀ al. (2025). Both sources satisfy their LRD selection criteria ( and ). The measured slope values are listed in TableĀ 2. In addition, both sources are unresolved in F444W, further supporting that they are typical LRDs.
@@
Altogether, we find that both sources are consistent with standard criteria for LRDs based on their morphology and āv-shapeā SED. In addition, both sources show the broad H line, making them strong Type I AGN candidates.
@@ 3 Results
3.1 Non-detections in both [Cāii] and 1.3 mm of two LRDs
We analyze the NOEMA data cubes and extract spectra at the positions of our sources using an aperture consistent with the beam size. The resulting spectra, shown in Fig.Ā 3, reveal no significant signal at the expected [Cāii]Ā frequency (). Assuming a [Cāii] line width of 250 km s-1 ā a typical value for high-redshift galaxies (BĆ©thermin etĀ al., 2020; Endsley etĀ al., 2022) ā we construct the velocity-integrated intensity map (moment-0). Fig.Ā 3 overlays the contours from the [Cāii]Ā moment-0 and 1.3 mm maps on the JWST RGB image, showing no detections () at the source positions and in the surrounding regions. We therefore conclude that neither the [Cāii] emission line nor the 1.3 mm dust continuum emission are detected for the two LRDs in the current NOEMA data.
We then determine 3 upper limits for both [Cāii]Ā and dust continuum. Using the rms of the 1.3 mm map (see Sect.Ā 2.2) and assuming our sources are not resolved, we derive 3 upper limits of 91 Jy and 151 Jy for ID9094 and ID2756, respectively. In addition, the 3 upper limits for [Cāii]Ā are calculated as three times the pixel-by-pixel rms of the moment-0 map, which gives upper limits of 0.157 Jy km s-1 and 0.310 Jy km s-1 for ID9094 and ID2756, respectively. The corresponding 3 upper limits on the [Cāii]ā emission-line luminosity are log( for ID9094 and log( for ID2756. Using , we further calculate the 3 upper limits on the molecular gas mass (Zanella etĀ al., 2018) to be log( for ID9094 and log( for ID2756.
Intriguingly, when requesting the NOEMA observations, we assumed these two galaxies to be DSFGs. Given their stellar masses from Xiao etĀ al. (2024), we estimated their total star formation rates (SFRtot) assuming that they lie on the star-forming main sequence (SFMS; Schreiber etĀ al., 2015) and predicted their expected [Cāii]Ā fluxes using the āSFRtot relation for SFGs (Schaerer etĀ al., 2020). Similarly, we predicted their 1.3 mm continuum fluxes by assuming an infrared SED shape consistent with typical dust-obscured galaxies (optically dark/faint galaxies; Xiao etĀ al., 2023). Given the depth of our NOEMA observations, we expected ID9094 to be detected at in [Cāii]Ā and in the 1.3 mm continuum, while ID2756 should be detected in [Cāii]Ā at . However, the complete absence of [Cāii]Ā and 1.3 mm emission suggests that these two LRDs are not the typical dusty SFGs initially assumed. This result aligns with our findings in Sect.Ā 2.4, where both sources exhibit a broad H emission line, being Type-I AGN candidates.
@@ 3.2 The hints of AGNs in two LRDs from FIR non-detections
In this section, we utilize the 3 upper limits on the [Cāii] emission line and the 1.3 mm continuum to investigate the possible presence of AGN activity in the LRDs. While both sources have already been identified as strong Type-I AGN candidates based on their broad H emission lines, our primary goal here is to investigate whether the FIR non-detections can effectively rule out the DSFG scenario for LRDs, independently of spectral information. Here, we focus primarily on ID9094, which benefits from deeper NOEMA observations compared to ID2756, allowing for more robust constraints. We assess the AGN contribution through two independent approaches: 1) UV-to-FIR spectral energy distribution (SED) fitting analysis and 2) the āSFRtot empirical relation.
@@ 3.2.1 SED fitting analysis
We perform a UV-to-FIR SED fitting analysis using CIGALE (Boquien etĀ al., 2019), which is based on an energy balance principle, fixing the redshift to . We assume a constant star formation history (SFH) and adopt Bruzual & Charlot (2003) stellar population models with the Calzetti etĀ al. (2000) dust attenuation law. We adopt a broad metallicity range of 0.004 to 0.02, dust attenuation to the rest-frame V band () values from 0 to 6 magnitudes, and ionization parameters log of -2.0. The AGN component is modeled using the Fritz etĀ al. (2006) template, with the AGN fraction (), i.e., the contribution of the AGN to the total IR luminosity, ranging from 0.01 to 0.8. The angle between the equatorial axis and the viewing angle at the line of sight () ranges from 40 to 90. With a fixed open angle of the AGN torus at 100 degrees, this range of refers to Type I AGN. We note that recent MIR observations (Wang etĀ al., 2025; Williams etĀ al., 2024; Barro etĀ al., 2024) suggest a lack of hot torus dust emission in LRDs, which may deviate from the Fritz etĀ al. (2006) template adopted. However, with the absence of deep MIR observations, we anticipate little change in the quality of SED fitting on our sample by only changing the AGN torus emission in the templates. On the contrary, the inclusion of continuum from accretion disk by limiting to Type I AGN is critical, as we robustly detected broad lines in H.
@@
Given that we only have upper limits in the FIR, we adopt a simple FIR dust model from Casey (2012), using the default dust emissivity index () and MIR power-law slope (). Compared to , which is adopted in some studies of dusty star-forming galaxies (e.g., Hildebrand, 1983; KovÔcs et al., 2006; Gordon et al., 2010; Gómez-Guijarro et al., 2022; Xiao et al., 2023), our choice of is more conservative and leads to slightly lower inferred dust temperatures for a given FIR slope. During the fitting process, we find that when the dust temperature () is allowed to vary across a defined range, the best-fit is strongly influenced by the upper boundary of the input range. This effect arises due to the absence of deep MIR data, which prevents us from constraining the maximum . Consequently, we cannot determine a precise best-fit value for , but the deep NOEMA data allow us to place a lower limit.
@@
To refine our constraints, we adopt an optimized fitting approach that allows us to constrain the minimum allowed using the NOEMA 1.3 mm 3 upper limits. In this approach, we fix during the fitting and perform multiple fits with different values, ranging from 30 K to 150 K. In addition, we also correct the effect of the cosmic microwave background (CMB) as an observing background on photometry, following da Cunha etĀ al. (2013). The CMB has the effect of raising the apparent dust temperature with increasing redshift, as the CMB temperature scales with redshift and provides an additional background source of radiation for dust heating. We assume , consistent with previous studies for galaxies (Sommovigo etĀ al., 2022). In this way, we derive the intrinsic FIR photometry starting from MIPS 70m, which is used for SED fitting. The best-fit SEDs, with and without AGN components considered, are presented in Fig.Ā 4.
@@
For ID9094, without including an AGN component (Fig.4-), we present three representative examples of best-fit SEDs with () K, () K, and () K. These correspond to results that the predicted flux density at 1.3 mm exceeds, matches, and falls below the observed 3 upper limits, respectively, under the energy balance assumption of CIGALE. At K ā a typical value for star-forming UV-bright galaxies in the REBELS survey at (Sommovigo etĀ al., 2022) ā the predicted 1.3 mm flux exceeds the observed 3 upper limit by nearly a factor of three. This discrepancy suggests that, with our current NOEMA observations, we should have detected the source at a significance level of , which is clearly not the case. As increases, the peak of the FIR SED moves towards the shorter wavelength and the fit improves. At K (Fig.4-), the predicted 1.3 mm flux aligns with the observed 3 upper limit. However, given that the actual signal-to-noise ratio at 1.3 mm is well below 3, this implies that the real would be greater than 110 K.
@@
Therefore, without including an AGN component, we obtain a lower limit of K. At fixed K, as in TableĀ 3, the best-fit IR luminosity is , yr-1 (we use this for the analysis in Sect.Ā 3.2.2), and stellar mass of . Remarkably, even this lower limit is significantly higher than the typical K observed in REBELS galaxies at similar redshifts (Sommovigo etĀ al., 2022). Similar findings have been reported in studies of 675 LRDs at , regardless of heating mechanism, whether AGN or star formation, where an average luminosity-weighted dust temperature of K is reported (Casey etĀ al., 2024). Such an extreme dust temperature is challenging to explain solely through heating by star formation. In typical star-forming galaxies, the dust temperature is primarily driven by the ultraviolet radiation field of young stars, which is usually insufficient to reach the high values observed here (see e.g., Harshan etĀ al., 2024; Laporte etĀ al., 2017).
@@
Based on the best-fit without AGN at fixed K, we derived the dust emitting region size pc, based on a black-body assumption777 and is the StefanāBoltzmann constant.. A crucial parameter to consider is the stellar surface density, given by pc-2 (similar findings are reported in, e.g., Baggen etĀ al., 2024). This is about one order of magnitude higher than the densest systems in the local universe (Hopkins etĀ al., 2010) and SFGs at observed by JWST (Morishita etĀ al., 2024; Schaerer etĀ al., 2024), highlighting a general issue for LRD models that rely solely on stellar components. Expanding the star formation region would reduce , but it would also decrease below the observed values (e.g., De Rossi etĀ al., 2018), leading to an inconsistency with our findings. These values are difficult to explain through star formation alone under standard assumptions and may suggest the contribution of additional heating sources, such as AGN activity. However, we emphasize that this interpretation remains model-dependent, and we further discuss such caveats in Sect.Ā 3.2.3.
| ID9094 | fixed [K] | log ] | SFR [yr-1] | [] |
| Without AGN | 110 | |||
| With AGN | 60 | |||
| With AGN | 40 |
@@
When the AGN component is included in the SED fitting (Fig.Ā 4-), the results become more physically plausible. Similar to Fig.4-, we also show three representative examples of best-fit SEDs with () K, () K, and () K. However, due to the degeneracy between the dust temperature and the AGN fraction, the current upper limit at 1.3 mm does not allow us to place a meaningful lower limit on . The observed photometry is consistent with SED fits with K and even marginally consistent with down to 40 K, a value comparable to those observed in AGN and/or quasar host galaxies at ( K; e.g., Walter etĀ al., 2022; Tsukui etĀ al., 2023; Tripodi etĀ al., 2023; Decarli etĀ al., 2023; Meyer etĀ al., 2025). At fixed K, the best-fit AGN fraction is , indicating a scenario where both AGN and star formation contribute significantly to the FIR emission. The corresponding best-fit , , and are listed in TableĀ 3. We also note that both sources are covered by the MEOW survey (JWST Cycle 3 GO-5407; PI: G. Leung; priv. comm.), which detects both in F1000W and ID9094 also in F2100W. Although the fluxes are not yet publicly released, preliminary analysis suggests that ID9094 exhibits significant hot dust emission in the MIR (G. Leung, priv. comm.).
@@
For ID2756, we perform SED fitting using the same strategy as for ID9094 (see AppendixĀ A). However, the current 1.3 mm observations for ID2756 (rms level 50.3 Jy beam-1) are not sufficiently deep to constrain the SED effectively. The lack of sensitivity prevents us from distinguishing between the DSFG and AGN scenarios based solely on the SED fitting results. To achieve meaningful constraints, deeper FIR/submillimeter observations are required. Specifically, observations at least as deep as for ID9094 (rms level 30.4 Jy beam-1) are necessary to provide tighter limits on the 1.3 mm flux density and help refine the SED fitting.
3.2.2 āSFRtot relation